A sample of a gas has a volume of 426mL at 330 K. What temperature is necessary for the gas to have a volume of 475 mL? Remember to convert mL to L. Round to the nearest whole number.
1
Expert's answer
2020-03-31T12:40:00-0400

V1 = 426 mL = 0.426 L

V2 = 475 mL = 0.475 L

T1 = 330 K

T2 = unknown


Solution:

Apply Charles' Law:

V1/T1 = V2/T2

Therefore,

T2 = V2*T1 / V1 = (0.475 L)(330 K)/(0.426 L) = 368 K

T2 = 368 K.


Answer: T2 = 368 K.